I recently posted this question and a tutor was kind enough to assist me. I have a question though. The answer the tutor came up with while assisting me with this question was not one of the available answers on my homework assignment. Below is the question and the answers. The tutor that assisted me came up with $141.98.

Find the monthly payments on a $2400 loan made for 18 months at 12% interest rate.


A. $146.36
B. $149.33
C. $288.00
D. $134.66

Do I need to contact my school that their answers are wrong?

Answer by Edwin McCravy(20054)

I recently posted this question and a tutor was kind enough to assist me. I
have a question though. The answer the tutor came up with while assisting me
with this question was not one of the available answers on my homework
assignment. Below is the question and the answers. The tutor that assisted me
came up with $141.98. 

Find the monthly payments on a $2400 loan made for 18 months at 12% interest
rate. 

A. $146.36
B. $149.33
C. $288.00
D. $134.66 

The monthly payment can be found by using the following formula:

  P = (Pv*R) / [1 - (1 + R)^(-n)] 
 
where

  Pv  = Present Value (beginning value or amount of loan) = $2400
  APR = Annual Percentage Rate (one year time period) = .12
  R   = Periodic Interest Rate = APR/ # of interest periods per year
      = .12/12 = .01
  P   = Monthly Payment
  n   = # of interest periods for overall time period (i.e., interest
        periods per year * number of years) = (12)(1.5) = 18

  P = (Pv*R) / [1 - (1 + R)^(-n)]
  P = (2400*.01)/ [1 - (1 + .01)^(-18)] 
  P = 146.3569149

rounded to the nearest cent:

  P = $146.36 which is your choice A.

Edwin